OpenStudy (anonymous):

What are the special products that appear in multiplying polynomials, and how are they recognized

OpenStudy (anonymous):

Special Products Product of a Sum: (a + b)^2 = a^2 + 2ab + b^2 Product of a Difference: (a - b)^2 = a^2 - 2ab + b^2 Product of a Sum & a Difference: (a + b)(a - b) = a^2 - b^2 Polynomials can be identified as special products by examining the characteristics of their terms.